### v1

# Quantum-private distributed sensing

Abstract
Quantum networks will provide unconditional security for communication, computation and distributed sensing tasks. We report on an experimental demonstration of private parameter estimation, which allows a global phase to be evaluated without revealing the constituent local phase values. This is achieved by sharing a Greenberger-Horne-Zeilinger (GHZ) state among three users who first verify the shared state before performing the sensing task. We implement the verification protocol, based on stabilizer measurements, and measure an average failure rate of $0.038\pm 0.005$ which we use to establish the security and privacy parameters. We validate the privacy conditions established by the protocol by evaluating the quantum Fisher information of the experimentally prepared GHZ states.
I Introduction
Quantum networks allow connected nodes to perform tasks such as multi-user quantum cryptography [1, 2], distributed quantum computation [3, 4], and distributed quantum sensing [5, 6, 7, 8]. The increased connectivity and sharing of resources over networks will likely prompt privacy-enabled protocols to protect sensitive information. A well known example of a private computing protocol is blind quantum computing, which allows clients to access remote quantum processors without revealing the computational algorithm to the servers [9, 3, 4]. Another established application is anonymous quantum communication, where users share cryptographic keys within a quantum network without revealing the identities of the keyholders to an eavesdropper, or other users within the network [10, 11, 12, 13, 14, 15]. Similar ideas from quantum cryptography can be incorporated into remote sensing tasks operating over quantum networks to provide both security and privacy features [16, 17, 18, 19].
Distributed quantum sensing typically involves spatially separated quantum sensors to monitor a global property, i.e., non-localised phenomena, such as magnetic fields, temperature sensing, or drifts between remote clocks [6]. The goal is to estimate a global value $\phi$ which is a linear function of local sensor values $\theta_{i}$, such as the weighted sum—this differs from multi-parameter estimation where each $\theta_{i}$ is estimated, for which it is often optimal to use local strategies [20, 21]. In certain scenarios, distributed sensing can exploit entanglement generation in networks to estimate the global parameter directly with Heisenberg-limited precision scaling [16, 22, 8]. Additionally, entanglement can be used to guarantee security, ensuring only authorised users learn the estimated parameters, and privacy—delegating the sensing task to an untrusted node to perform measurements without learning the parameters [17, 18]. Recently, a two-user scheme has been proposed [23] and demonstrated [24] by sharing a Bell pair between two parties, one who performs the sensing task while only the other learns the estimated parameter.
Figure 1: Sensor nodes in a network monitor a global function of parameters while local values $\{\theta_{i}\}$ remain secret. An honest verifier establishes private pairwise channels with each node (not shown). (a) An untrusted server sends $N_{\textrm{t}}$ copies of the resource state to all nodes, who store them locally in a quantum memory. They perform the verification protocol to ensure they share a state close to a GHZ state. (b) If verification is successful, the nodes encode their local parameters on one copy of the shared state to perform parameter estimation.
Here we consider multi-user quantum-private distributed sensing, a task known as private parameter estimation (PPE), for which security bounds are recently established in Ref [25]. We illustrate the conceptual arrangement of the PPE task in Fig.1. The $n$ remote sensors are used to estimate a global phase, $\phi$, which is a sum $\phi\doteq\sum_{i}\theta_{i}$ of $\theta_{i}$ local phase values, with $i\in\{1,\dots,n\}$. The goal is to prevent information about the local $\theta_{i}$ from being learned by any node or eavesdroppers monitoring the communication. Analogous to quantum conference key agreement [26, 27, 15, 28], the correlations of the $n$-partite GHZ state, |$\ket{GHZ}\doteq(\ket{0}^{\otimes{n}}+\ket{1}^{\otimes{n}})/\sqrt{2}$ where |$\ket{0}$ and |$\ket{1}$ are computational basis states, are exploited to perform the distributed sensing task privately. The nodes must first verify that they are sharing the GHZ state to ensure they are performing the protocol in private and with integrity. Once this is achieved, the $\theta_{i}$ are encoded onto the shared GHZ state by coupling each node’s qubit to their respective quantum sensor, then each node measures their qubit and announces the result.
We now outline the state verification protocol [25]. One of the $n$ nodes is designated as the verifier, who is assumed to be honest and has pairwise private channels with each remaining node. An untrusted quantum server prepares and distributes $N_{t}$ copies of the $n$-partite GHZ state to all nodes who store them locally in quantum memories. The verifier randomly chooses half of the copies ($N_{t}/2$) and instructs the group to measure them according to one of the stabilizers, $K_{i}$. The set of stabilizers are given by cyclic permutations of $K_{i}=-X^{(1)}X^{(2)}\dots Y^{(i)}Y^{(i+1)}\dots X^{(n-1)}X^{(n)}$ for $i=1,2,\dots,n$ and $K_{i=n+1}=X^{\otimes n}$. By definition, the measurement outcomes of $K_{i}$ result in a $+1$ eigenvalue for the $n$-qubit GHZ state. Each node performs their respective measurement then communicates their outcome to the verifier directly using their private channels. The verifier computes a failure rate, $f$, as the fraction of copies corresponding to a $-1$ eigenvalue. Provided $f$ is small, one of the remaining untested copies is chosen at random as the target copy, $\rho_{j}$, and is used for the parameter estimation. This verification scheme is closely related to the one in Ref. [29], which is generalised to verify graph states using the full set of stabilizer measurements requiring $2n$ measurements, while in Ref. [25] only $n+1$ measurements are required for the GHZ state. This verification protocol guarantees a lower bound on the fidelity $F(\rho_{j},\ket{GHZ})$ of the target copy with the GHZ state up to some nominal probability,
$\mathbb{P}\bigg{(}F(\rho_{j},\ket{GHZ})\geq 1-\frac{2\sqrt{c}}{n}-2nf\Bigg{)}\geq 1-n^{1-\frac{2mc}{3}}.$ (1)
$\mathbb{P}(\textbf{A})$ is the probability that A is true, given the failure rate $f$, number of users $n$, and two positive variables $c$ and $m$ which relate to the statistical certainty based on the total number of copies, $N_{t}=\lceil 2mn^{5}\text{log}(n)\rceil$, used in the protocol (more details in the Appendix 2). Minimising $c$ improves the fidelity lower bound but decreases the probability of the inequality holding true. The parameter $m$ can be increased commensurately to offset the reduction by $c$, but this requires more test copies. Use of quantum memories in the verification step safeguards against an adversary swapping out non-test copies with non-private states, $|+\rangle^{\otimes{n}}$ where $|+\rangle\doteq(|0\rangle+|1\rangle)/\sqrt{2}$ which allows the adversary to learn the local phases from the outcomes. We assume the future availability of such quantum memories and prepare multi-partite GHZ states in the lab to investigate the implementation of the verification scheme presented in Ref. [25].
If the verification protocol is successful, the verifier announces the target copy, $\rho_{j}$, to be used by the group to encode their local phase and perform the parameter estimation. The remaining $N_{t}/2-1$ untested copies are discarded [^1] Note that it might be possible to use some of these discarded copies in the protocol, however the security implications of this are not yet clear., which is needed to satisfy the security demands for the verification scheme [25]. The private sensing task proceeds with each node coupling their sensor to the target qubit, which is equivalent to applying local unitary encoding operators, $U(\theta_{i})=\exp(-i\theta_{i}\sigma_{Z}/2)$, where $\sigma_{Z}$ is the Pauli-Z matrix, on their qubit. Finally, each node measures in the $X$ basis and announces their outcome. The parity of the outcomes is used to estimate $\phi$ by repeating the protocol $\nu$ times, leading to a variance scaling of $\textrm{var}[\phi]\propto(\nu)^{-1}$ in the ideal case. To demonstrate the values of $\theta_{i}$ are private, we adopt the notions of privacy introduced in Ref. [25] based on the quantum Fisher information (QFI).
Figure 2: Experimental setup preparing a three-qubit GHZ state and implementing PPE. A Ti:sapph laser pumps two entangled photon-pair sources (EPS), one photon from each source is sent to a polarising beamsplitter gate (PBS-G). A motorised stage implements a tunable delay, $\delta\tau$. Each photon is sent to a measurement stage consisting of a quarter-wave plate (QWP), half-wave plate (HWP), and polarising beamsplitter (PBS) with both outputs monitored by superconducting nanowire single photon detectors (SNSPD). Local phases $\theta_{i}$ are implemented for the three nodes using a QWP-HWP-QWP. The fourth photon is measured as a trigger.
II Results
We experimentally prepare a three-qubit GHZ state to investigate the PPE protocol for $n=3$ users, see Fig 2. Two entangled photon-pair sources are implemented using aperiodically-poled KTP crystals, phase-matched for Type-2 parametric down conversion, embedded in a Sagnac loop [30]. Both sources are pumped by a mode-locked titanium:sapphire laser with a 80 MHz repetition rate, 1.3 ps pulse duration, and centred at 775 nm, to produce polarisation-entangled photon pairs at 1550 nm. One photon from each source is sent to a polarising beamsplitter (PBS) and each of the four photons are sent to polarisation analysers which perform tomographic measurements. The fourth photon in the setup is projected in the $X$ basis to obtain the GHZ state,
[…]
where $(1-p)$ is strength of added noise, which we use to test the protocol robustness, via a relative temporal delay $\delta\tau$ added to one photon in the PBS gate. We set $\delta\tau=0$ mm to maximise $p$ then perform full quantum state tomography (QST) to reconstruct the density matrix, $\rho$. At this setting we measured the three-qubit GHZ state fidelity as $0.923\pm 0.005$ and a state purity of $0.865\pm 0.009$. For a pure GHZ state both values are unity, however in practice several limitations exist; multi-photon events that arise from our probabilistic sources, mode mismatch on the PBS gate and finite polarisation extinction.
To implement the PPE protocol, the three nodes perform the verification protocol by measuring the stabilizers $K_{i}$, which we pick from a subset of the recorded QST measurements. For the optimal state we obtain an average failure rate $f=0.039\pm 0.005$ and from Ref [25] the verification step is successful if $f\leq 1/(2n^{2})$, in our three-node case this sets a threshold of $f\leq 0.05\dot{5}$ which our state satisfies. Using $f$ we establish a lower-bound fidelity of $0.769\pm 0.006$, via $F(\rho,GHZ)\geq 1-2nf$ when assuming infinite resource-state copies ($c\rightarrow 0$), which is smaller than the fidelity obtained via QST. For our three-node scenario it would be more optimal to use QST, however the measurements scale exponentially with $n$, while the verification protocol scales linearly requiring $n+1$ stabilizers. We also establish an upper-bounded privacy parameter, $\varepsilon_{p}\leq 1.3\pm 0.2$ using $\varepsilon_{p}\leq 24n^{-2}\sqrt{2nf}$, again assuming infinite copies, see Appendix 2 for details. For PPE, $\varepsilon_{p}=0$ corresponds to perfect privacy, wherein $\theta_{i}$ cannot be learned by any node nor an eavesdropper and $\phi$ can be estimated ideally [25]. Using the density matrix reconstructed from QST, we calculate the QFI and obtain $\varepsilon_{p}=0.005\pm 0.002$, which is much smaller than the upper bound from the verification protocol—we will follow up on this in the next section. We implement the phase sensing task for two global phases $\{\phi_{1},\phi_{2}\}$, see Fig. 3. To ensure each phase element encodes $\theta_{i}$ correctly we perform a full phase sweep for each, see Appendix 2 for details, then we independently set $\theta_{i}$ to prepare each $\phi$. To estimate $\phi$, all nodes measure in $X$ then compute the expectation value using all outcomes and evaluate $\phi=\textrm{cos}^{-1}[\langle X^{\otimes n}\rangle]$. The accuracy of the estimate $\hat{\phi}$ to the true value $\phi$ is evaluated as the distance, $\|\hat{\phi}-\phi\|$, see Fig. 3 (a), which approaches zero with a pure GHZ state and error-free encoding of $\theta_{i}$. We similarly evaluate the distance for $\theta_{i}$, but with only outcomes of the $i$-node; this approach can recover $\theta_{i}$ using the non-private state. The distances for estimating $\phi$ are over an order of magnitude lower than for $\theta_{i}$, and remain constant with increasing measurement samples $\nu$, which shows $\phi$ was successfully estimated. We then evaluate the variance in the estimates for the $\phi$ as a function of $\nu$, see Fig. 3 (b), to obtain the precision scaling. The experimentally measured variances generally approach the scaling of $\nu^{-1}$, which is attainable with a pure GHZ state and optimal measurements, which confirms our estimate of $\phi$ is near optimal. We also plot the precision scaling for $\theta_{i}$ using the directly evaluated privacy parameter, $\varepsilon_{p}=0.005\pm 0.002$. This represents the upper bound on the precision scaling for $\theta_{i}$, based on the state quality, and is only realisable through an optimal (but generally unknown) measurement.
Figure 3: Private parameter estimation for two global phases $\{\phi_{1},\phi_{2}\}$, all nodes measures in $\sigma_{X}$ to estimate $\phi$; this is repeated $\nu$ times. (a) The accuracy of the estimate $\hat{\phi}$ is evaluated as the distance for increasing $\nu$. Solid circles show experimental data points for $\phi$, unfilled square and diamonds represent the estimates for $\theta_{i}$. (b) The precision scaling is evaluated as the variance of estimate $\hat{\phi}$ with increasing $\nu$. Solid circles are experimental data, while solid lines show theoretically expected scaling. The purple curve scales as $\nu^{-1}$ is the best performance for $\phi$ using a pure GHZ state, while the red curve scales as $(\nu\varepsilon_{p})^{-1}$ is the best case for $\theta_{i}$.
We now investigate the verification protocol further by considering three main properties; the robustness to noisy GHZ states, comparing the privacy bounds with a direct evaluation of $\varepsilon_{p}$, and the impact of finite copy statistics. To test the robustness of the PPE task we prepare several GHZ states with added noise by setting the $\delta\tau$ away from 0 and compute $f$ in each case see Fig 4(a). For each state we perform QST to obtain fidelity values and we calculate the lower-bound fidelity using the stabilizers in Fig 4(b). In all cases, the lower bound is much smaller than the measured fidelity value, which ensures the protocol is robust to noise as the verification never overestimates the quality of the state.
Following the approach in Ref. [25] we will verify the privacy independently using the quantum Fisher information (QFI), a standard tool for benchmarking quantum metrology schemes, see Fig 4(c). The QFI quantifies the extractable information about a parameter encoded on a system with an optimal measurement. For a pure GHZ state and measuring in $X$ on all qubits, one can estimate $\phi$ with a theoretical QFI$=1$. Using the density matrices $\rho$ in our experiment we calculated the highest QFI=$0.90\pm 0.02$ when $\delta\tau=0$ while higher noise results in a lower value. We then evaluate the privacy parameter, $\varepsilon(\theta_{i})$, directly from $\rho$ for each local node value, $\theta_{i}$. We do this by calculating the reducible QFI of $\theta_{i}$, when encoded by the $i$-node and when the other $j\neq i$ nodes encode functions of $\theta_{i}$ such that the resulting QFI is minimised. For a GHZ state, the reducible QFI of any local parameter $\theta_{i}$ is zero, by letting the other nodes encode the local parameters $\theta_{j}=-\theta_{i}/(n-1)$, see Appendix 4 for details. For the non-private state $|+\rangle^{\otimes n}$, we obtain $\varepsilon_{p}=1/n^{2}$. We calculate $\varepsilon(\theta_{i})$ for all cases of $\delta\tau$, see Fig 4(c), we also evaluate the protocol privacy parameter, $\varepsilon_{p}\doteq\max_{i}\varepsilon(\theta_{i})$ for each case. The experimentally measured values for $\varepsilon_{p}$, of all tested $\delta\tau$, are smaller than the calculated value for the non-private state by over one order of magnitude—this confirms the privacy of the states we prepared.
[…]
So far we have been evaluating the PPE task in the asymptotic regime by assuming infinite copies of states, however in practice only a finite number will be available. At $\delta\tau=0$ we repeat the verification protocol to construct a histogram of the measured failure rates, see Fig 4(d). In total we recorded over $1.2\times 10^{7}$ detection events evenly across the stabilizer bases. We plot histograms of the measured failure rates and observe the spread of $f$ reduce as the number of test copies increases, as expected. It is necessary to optimise the number of copies devoted to testing while maintaining confidence on the parameter $f$, which is captured by the parameters $c$ and $m$, see Appendix 2 for details.
III Discussion
We demonstrated a quantum-private distributed sensing task involving three nodes that measure a global phase $\phi$ without revealing the local sensor values $\theta_{i}$—the precision in estimating $\phi$ is at least two orders of magnitude better than for any $\theta_{i}$. We assumed the availability of pairwise private channels between each node and the honest verifier for securely reporting outcomes in the verification protocol, this is a common requirement that is satisfied for privacy-enabled protocols [11, 10, 13, 15, 14]. Our work highlights a number of deficiencies in the implemented PPE protocol. First, we found the bounds on the fidelity of the GHZ state and privacy parameter of the task, which are obtained from the failure rate in the verification protocol, were exceedingly loose. Tighter security, for example by rethinking the privacy definition in a composable security framework, would make networked distributed sensing more resource efficient, and there is no fundamental reason preventing these from being derived in the future. Second, in its current form, the PPE protocol requires all copies of the shared states to be stored in memories before they can be used. The finite bandwidth even of multi-mode memories would strongly limit the number of copies that can be used, and as we have seen from the performance penalties of our finite round analysis, PPE would effectively become impractical. One may alleviate the need for large memories by introducing a trusted random variable that chooses whether a distributed state in a sequence is used or not for verification, such that each node only requires a single-qubit quantum memory [31]. Alternatively, one may consider GHZ-state certification protocols which exploit fast, low-loss optical switches to randomly select a copy out of an ensemble which can be used for parameter estimation, while the remaining copies are certified [32, 33, 34]. While the PPE protocol we demonstrated has some caveats, we also identified pathways for improvements, which, once addressed, will unlock a wider range and more efficient distributed sensing applications for quantum networks.
Acknowledgements
We thank A. Pickston, L. Stroh, D. Markham and N. Shettell for helpful discussions. This work was supported by the UK Engineering and Physical Sciences Research Council (Grant Nos. EP/T001011/1.). FG acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the DFG Individual Grant BR2159/6-1.
Author contributions
JH and JWW are co-first authors of this work. JWW and AF conceived the project. JH, JWW, RMJB performed the experiment including taking data and analysing the results. FG and EG developed the theoretical tools used in the analysis. All authors contributed to writing and revisions of the manuscript.
Competing interests
The authors declare no competing financial or non-financial interests.
Appendix
4 Private parameter estimation protocols
In the main body we used two protocols that were initially introduced in Ref. [25]. While these are separated as two protocols, we remark that the second protocol is only implemented if the first protocol does not fail. The first protocol, Verification, is used to determine how close the distributed state is to the ideal GHZ state. The second protocol, Parameter estimation, implements the private distributed sensing task in its entirety. These protocols are outlined here for completeness.
Protocol 1: Verification [29]
1: An untrusted source generates $N_{t}$ copies of the $n$-qubit resource state, for the group of $n$ nodes. For each copy, the $j$-th qubit is sent to the $j$-th node, where $j\in\{1,2,\dots,n\}$.
2: For $i\in\{1,2,\dots,n\}$, do the following: (a) The Verifier selects $N_{t}/(2n)$ copies independently and uniformly at random. (b) For each selected copy, the Verifier instructs each party to perform the measurement corresponding to their part of the stabilizer operator $K_{i}$. (c) For each selected copy, the parties send their measurement outcome to the Verifier, who computes the outcome of the measurement of $K_{i}$. The copy passes the test if the measurement outcome of $K_{i}$ is +1. Let $N_{\text{pass},i}$ be the number of copies that pass the stabilizer test for $K_{i}$.
3: The Verifier chooses a single copy from the remaining $N_{t}/2$ copies that were not used for verification, uniformly at random. The chosen single copy is called the target copy $\rho$. The others are discarded.
4: The average failure rate $f=1-2N_{\text{pass}}/N_{t}$ is calculated and if $f\leq\frac{1}{2n^{2}}$, the parties use the target copy $\rho$ for their task; otherwise, the verification of the GHZ state failed and the target copy is discarded.
Table 1: Verification protocol. The Verifier is assumed to be honest. Input: The verifier requests $N_{t}$, the total number of copies of the resource state that are distributed. The set of stabilisers $\{K_{i}\}_{i=1}^{n}$ of the GHZ state is The set of stabilisers, are given by cyclic permutations of $K_{i}=-X^{(1)}X^{(2)}\dots Y^{(i)}Y^{(i+1)}\dots X^{(n-1)}X^{(n)}$ for $i=1,2,\dots,n$ and $K_{i=n+1}=X^{\otimes n}$ [35]. Output: a target copy close to the GHZ state or abort.
Protocol 2: Parameter estimation [25]
1: Repeat the following $\nu$ times. (a) The nodes run the Verification protocol, if the computed failure rate $f>\frac{1}{2n^{2}}$ then the protocol fails and the users abort. If Verification does not fail, the nodes share a resource state close to a GHZ state. (b) Each node $j$ encodes their local parameter $\theta_{j}$ by applying $U_{j}=e^{-i\theta_{j}\sigma_{Z}/2}$ on their qubit. (c) Each node measures their qubit in the $X$ basis and announces the measurement outcome.
2: The outcomes are used to compute the expectation value, $\langle X^{\otimes n}\rangle$, and the global phase is evaluated as $\hat{\phi}=\textrm{cos}^{-1}[\langle X^{\otimes n}\rangle]$ in each iteration. Conditional on successful verification rounds, the achievable precision scaling with respect to these rounds approaches $\nu^{-1}$ in the ideal case.
Table 2: Secure networking sensing protocol [25]. Input: The parties choose a level of precision for estimating $\phi$ which sets the number of rounds $\nu$. Each node encodes local phases $\theta_{1},\ldots,\theta_{n}$. Output: Estimation of $\phi=\sum_{i}\theta_{i}$, with $\varepsilon_{i}$-integrity and $\varepsilon_{p}$-privacy.
5 Analysis of privacy and integrity
Privacy
From the parameters defined by the verification model presented in Ref. [25] one can construct a bound for the privacy that is captured by the protocol and provides an upper bound on $\varepsilon_{p}$ given as follows:
$\varepsilon_{p}\leq\frac{24}{n^{2}}\sqrt{\frac{2\sqrt{c}}{n}+2nf},$ (SM1)
where $n$ is the number of sensors and $f$ is the failure rate of the stabilizer tests, averaged over all the $n$ tested stabilizers; conditional on the success from the verification protocol we note that $f\leq 1/(2n^{2})$. The positive constant $c$ is a free variable constrained by $43/2m<c<(n-1)^{2}/$ 4. The verifier can choose the failure rate $f$ and the free variable $c$ to obtain a satisfactory level of privacy and integrity in their protocol and decrease $\varepsilon_{p}$ to an arbitrarily small value.
In particular, $\varepsilon_{p}$-privacy quantifies the local sensor information leakage to an adversary. We extend the privacy definition from [25] to the case of $n$ parameters encoded on a multipartite state via unitary encoding.
Definition 1.
Let $\rho_{\theta_{1},\dots,\theta_{n}}$ be a $n$-partite state encoded with $n$ parameters through the local unitaries $U_{j}=\exp(-i\theta_{j}H_{j})$, where $H_{j}$ is a Hermitian operator on the $j$-th system. Then, the state is said to be $\varepsilon_{p}$-private if, $\forall\,j\in[1,n]$, there exists choices for the other local phases, $\phi_{k}(\theta_{j})\mbox{ for }k\neq j$, such that:
$\quad F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})\leq\varepsilon_{p}(h_{\max}-h_{\min})^{2},$ (SM2)
where $F_{Q|\theta_{j}}$ is the QFI calculated with respect to the parameter $\theta_{j}$ and where $h_{\max}$ ($h_{\min}$) is the largest (smallest) eigenvalue of the operator $\sum_{j=1}^{n}H_{j}$.
Note that we expect $\varepsilon_{p}\in[0,1]$ since for the trivial choice $\phi_{k}=\theta_{j}$ we get $F_{Q|\theta_{j}}(\rho_{\theta_{j},\dots,\theta_{j}})\leq(h_{\max}-h_{\min})^{2}$ (this result follows from Theorem 1 in [36]).
For the unitary encoding operated by the nodes in our experiment, we have $H_{j}=\sigma_{Z}/2$, which corresponds to the Definition 2 which recovers the definition provided in Ref. [25].
Definition 2.
The $n$-partite state $\rho_{\theta_{1},\dots,\theta_{n}}$, encoded by the local unitaries $U_{j}=\exp(-i\theta_{j}\sigma_{Z}/2)$, is said to be $\varepsilon_{p}$-private if, $\forall\,j\in[1,n]$, there exists $\phi_{k}(\theta_{j})\mbox{ for }k\neq j$ such that:
$\quad F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})\leq\varepsilon_{p}n^{2}.$ (SM3)
Integrity
We now discuss the $\varepsilon_{i}$-integrity of the protocol. The authors of Ref. [25] quantify the deviation in accuracy and in precision of the estimator of the protocol, when the shared resource deviates from a GHZ state. As such, these measures depend on an upper bound on the average trace distance between the encoded resource state and the encoded GHZ state –averaged over the parameter estimation rounds [25]. Thus, we have the following definition.
Definition 3.
Let $\tau_{j}$ be the target state of the $j$-th verification protocol after being encoded with the local phases, and let $T(\rho,\sigma)$ be the trace distance between $\rho$ and $\sigma$. Then, the metrology protocol is $\varepsilon_{i}$-integrous if:
$\frac{1}{\nu}\sum_{j=1}^{\nu}T(\tau_{j},\tau_{\phi})\leq\varepsilon_{i},$ (SM4)
where $\tau_{\phi}$ is the encoded GHZ state,
$\tau_{\phi}=\ket{\Phi}\bra{\Phi},\quad\ket{\Phi}=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes n}+e^{i\phi}|1\rangle^{\otimes n}).$ (SM5)
The deviations in accuracy and precision derived in [25] are obtained for a specific estimator $\hat{\phi}$ of the global phase. In particular, let $\tau_{\phi}$ be the ideal encoded state, (SM5). Then, the expectation value of $\sigma_{X}^{\otimes n}$ on $\tau_{\phi}$ is:
$g(\phi)=\text{Tr}[\sigma_{X}^{\otimes n}\tau_{\phi}]=\cos\phi.$ (SM6)
On the other hand, the expectation value of $\sigma_{X}^{\otimes n}$ is approximated by the average of its measurement outcomes over the $\nu$ metrology rounds, when $\nu\gg 1$ and when the resource state is close to the GHZ state:
$\hat{g}=\frac{1}{\nu}\sum_{j=1}^{\nu}m_{j},$ (SM7)
where $m_{j}$ is the outcome of $\sigma_{X}^{\otimes n}$ in the $j$-th metrology round (i.e., computed on $\tau_{j}$). The estimator employed in [25] is defined as:
$\hat{\phi}$ $=g^{-1}(\hat{g})$
$=\arccos(\hat{g}).$ (SM8)
The authors in [25] argue that the estimator in (SM8), for $\nu\gg 1$, is unbiased when the resource state coincides with the ideal state ($\tau_{j}=\tau_{\phi}$ for all $j$). Then, Theorem 1 in [25] provides a bound on the bias of the estimator computed on the actual resource state:
$|\mathbb{E}(\hat{\phi})-\phi|\leq\frac{2o\varepsilon_{i}}{|\sin\phi|},$ (SM9)
where $\mathbb{E}(\hat{\phi})$ is the expectation value of (SM8) on the actual resource state, while $o$ is the maximum magnitude of the eigenvalues of the observable which is being measured for parameter estimation; in our case $o=1$.
Similarly, one can argue that the variance of the estimator in (SM8), assuming that the resource state is the same in every parameter estimation round ($\tau_{j}=\tau$), can be computed as [25]:
$\Delta^{2}\hat{\phi}=\frac{1-(\text{Tr}[\sigma_{X}^{\otimes n}\tau])^{2}}{\nu\,\sin^{2}\phi}.$ (SM10)
Now, when the resource state is the encoded GHZ state, the variance simplifies to $1/\nu$. Theorem 1 in [25] provides a bound on the deviation of the precision of the estimator when the resource state is not ideal:
$\left\lvert\Delta^{2}\hat{\phi}-\frac{1}{\nu}\right\rvert\leq\frac{4\varepsilon_{i}(2\nu^{-1}+\varepsilon_{i})}{\sin^{2}\phi}.$ (SM11)
We evaluate the bounds in (SM11) and (SM9) by replacing $\mathbb{E}(\hat{\phi})$ with $\hat{\phi}$ and $\Delta^{2}\hat{\phi}$ with the formula in (SM10), where we choose $\tau=\sum_{j}\tau_{j}/\nu$:
$\Delta^{2}\hat{\phi}$ $=\frac{1-(\mathbb{E}(\hat{g}))^{2}}{\nu\,\sin^{2}\phi}$ (SM12)
$\approx\frac{1-(\hat{g})^{2}}{\nu\,\sin^{2}\phi}.$ (SM13)
Experimental validation of integrity
The $\varepsilon_{i}$ parameter is the average trace distance between the stored state after verification and an ideal state. This can be calculated through one of two ways, either taken directly from the experiment or taken from the current theoretical framework outlined in [25]. We will show that the experimentally obtained $\varepsilon_{i}$ term leads to a tighter bound than the current theoretical term. First, using the experimental fidelity achieved, the upper bound on the trace distance via the fidelity is $\textrm{Tr}(\rho-\sigma)\leq\sqrt{1-F(\rho,\sigma)}$, where $\varepsilon_{i,\textrm{exp}}=\sqrt{1-F(\rho,\sigma)}$. The second method relies exclusively on the information provided by the protocol in [25], which states that after the verification protocol the fidelity must be above a certain threshold— Eq. (1) in the main text. Using this lower bound and the trace distance via the fidelity, we get $\varepsilon_{i,\textrm{theo}}=\sqrt{2\sqrt{c}/n+2nf}$. We can now compare these two integrity metrics, using one of the phase sweep plots presented in Fig. SM2. The accuracy and precision and hence the integrity of the sensors can be plotted, as presented in Fig.SM1. The accuracy is indeed loose in both regards of $\varepsilon_{i,\textrm{exp}}$ and $\varepsilon_{i,\textrm{theo}}$, however our experimentally obtained integrity parameter is substantially tighter than the current theoretical framework. The measured precision appears to be up to seven orders of magnitude away from the theoretical bound, in both the $\varepsilon_{i,\textrm{exp}}$ and $\varepsilon_{i,\textrm{theo}}$ expressions. It should be noted that for $\varepsilon_{i,\textrm{exp}}\rightarrow 0$, the theoretical expression scales exponentially towards our precision. This hints that modelling $\varepsilon_{i,\textrm{exp}}$ through the trace distance measure leads to a loose criteria, but not as loose as previously established through $\varepsilon_{i,\textrm{theo}}$, which does not get close to the characteristics determined from the experimental data—thus concluding and quantifying the looseness. We have therefore provided an alternative, tighter, bound in both the accuracy and precision of the protocol, leading to the integrity parameter being more honest.
Figure SM1: The accuracy and precision of the protocol, defined with respect to $\varepsilon_{i}$. In both plots, the green line is defined w.r.t $\varepsilon_{i,\textrm{theo}}=\sqrt{2\sqrt{c}/n+2nf}$ and the purple line is defined w.r.t $\varepsilon_{i,\textrm{exp}}=\sqrt{1-F(\rho,\sigma)}$. (a) shows the accuracy, as in (SM9), of the measured global phase (black line). Specifically, the green and purple are the right hand side of the inequality and the black line is the left hand side of the inequality in (SM9). (b) shows the precision, a more direct measure of integrity as in (SM11) it is w.r.t a certain number of parameter estimation rounds $\nu$. The green and purple lines are the right hand side and the black line is the left hand side of the inequality in (SM11). This is taken at $\nu=3200$, $c=0.25$, with $f=0.047$ as measured in (SM9), and the fidelity used is 0.905, the lower bound of 0.907(2).
Figure SM2: Measured visibilities of each node phase encoding stage. In each case, two phase shifters are fixed in phase, while one is rotated. The measured visibilities are $91.5(3)\%$, $92.0(3)\%$ and $91.9(3)\%$ for (a), (b) and (c) respectively. For completeness we note that the fourth qubit is also projected in X in the setup, hence the fourth qubit label.
6 Supplementary Discussion
We have alluded to the fact that the protocol is rather pessimistic in the lower bound of the fidelity of the resource state which leads to large bounds on the privacy and integrity figures of merit. To highlight the distance between the lower bound fidelity and the actual fidelity we plot in FIG SM3 the failure rate as a function of fidelity in the asymptotic limit of infinite copies. The solid black line is the estimated fidelity using protocol 1, which intersects the maximum failure rate (red dashed line) at 0.67 fidelity. However, the experimental resource state with a similar average failure rate of 0.0525 has an estimated fidelity of 0.89, the horizontal distance along the red dashed line represents the under performance of the protocol at estimating the fidelity of this state, which is 22% under estimate. We can consider what would happen if we now lower the maximum failure rate threshold since this is a free parameter by the verifier. If we set $f$ to 0.0183 then we will in fact accurately estimate the fidelity of the experimental resource state to be 0.89. However, we have observed the in FIG SM4 that the probability of a resource state to pass the stabiliser measurements follows a normal distribution and if the average failure rate of our resource state was 0.047, then the probability of one of these states passing the verification protocol with $f=0.0183$ would be around 0.2%.
Figure SM3: Failure rate is plotted against fidelity. The green line models the asymptotic limit of the verification protocol with the addition of dephasing noise and blue for depolarising noise. The black line is the asymptotic limit considering no noise model. The blue dots are the experimentally obtained fidelities and their corresponding failure rates, calculated by full quantum state tomography.
Figure SM4: Histogram of errors in measured verification rounds. Around 1068 rounds of each respective stabiliser are obtained, totalling 3204 measurements dedicated to the verification stage. This is repeated 3200 times, leading to a histogram whose mean value is $f=0.047$. Of these measurements, 3127 are within the error threshold of 0.055.
7 Direct calculation of QFI for privacy statement
In this section we perform a direct calculation of the QFI appearing in the privacy definition that applies to our experiment, Definition 2. This is achieved by employing the tomographically-reconstructed resource state of our experiment. The goal is to compare the direct calculation of the QFI, which would lead to a tight privacy bound $\varepsilon_{p}$, with the upper bound on $\varepsilon_{p}$ provided by [25] and reported in (SM1).
We start by recalling that the QFI of a state $\rho_{\theta}$, obtained by encoding $\theta$ into $\rho$, is given by [37]:
$F_{Q|\theta}(\rho_{\theta})=2\sum_{k,l}\frac{\left\lvert\braket{\psi_{k}}{\partial_{\theta}\rho_{\theta}}{\psi_{l}}\right\rvert^{2}}{\lambda_{k}+\lambda_{l}},$ (SM14)
where the sum only runs over indices for which $\lambda_{k}+\lambda_{l}\neq 0$ and where $\lambda_{k}$ and |$\ket{\psi_{k}}$ are obtained from the spectral decomposition of $\rho_{\theta}$: $\rho_{\theta}=\sum_{k}\lambda_{k}\ket{\psi_{k}}\bra{\psi_{k}}$ | (where $\lambda_{k}$ can also be zero). When the parameter $\theta$ is encoded by the unitary transformation:
$\rho_{\theta}=e^{-i\theta A}\rho e^{i\theta A},$ (SM15)
one can show that the QFI reduces to:
$F_{Q|\theta}(\rho_{\theta})=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{\lambda_{k}+\lambda_{l}}\left\lvert\braket{\varphi_{k}}{A}{\varphi_{l}}\right\rvert^{2},$ (SM16)
where now $\lambda_{k}$ and |$\ket{\varphi_{k}}$ are obtained from the spectral decomposition of the unencoded state $\rho$: $\rho=\sum_{k}\lambda_{k}\ket{\varphi_{k}}\bra{\varphi_{k}}$ |, implying that the QFI in (SM16) is independent of the value of $\theta$. Moreover, if the state $\rho$ is pure, $\rho=\ket{\varphi}\bra{\varphi}$ |, it holds:
$F_{Q|\theta}(\rho_{\theta})=4\left[\braket{\varphi}{A^{2}}{\varphi}-\left(\braket{\varphi}{A}{\varphi}\right)^{2}\right].$ (SM17)
In order to determine the privacy parameter $\varepsilon_{p}$, let us fix for the moment $j=1$ in (SM3), implying that we want to make a statement about the privacy of the local phase $\theta_{1}$ when encoded in our resource state $\rho$. In order to find the best possible privacy parameter $\varepsilon_{1}$, we solve the following optimization problem:
$\alpha_{1}=\min_{\phi_{k}(\theta),\,k\neq 1}F_{Q|\theta_{1}}(\rho_{\theta_{1},\phi_{2}(\theta_{1}),\dots,\phi_{n}(\theta_{1})}),$ (SM18)
and then choose $\varepsilon_{1}=\alpha_{1}/n^{2}$. The optimization is done over all possible choices of functions $\phi_{k}(\theta)$, such that the resulting state is functionally independent of $\theta_{1}$ –or as independent as possible. If we repeat this procedure for each $j$, we can then set $\varepsilon=\max_{j}\varepsilon_{j}$ and claim that the resource state $\rho$ is $\varepsilon_{p}$-private according to definition (SM3). To solve the optimization in (SM18), we compute the QFI, with respect to $\theta_{1}$, of the encoded resource state:
$\rho_{\theta_{1},\phi_{2}(\theta_{1}),\dots,\phi_{n}(\theta_{1})}=e^{-i\theta_{1}\frac{Z_{1}}{2}}e^{-i\phi_{2}(\theta_{1})\frac{Z_{2}}{2}}\cdots e^{-i\phi_{n}(\theta_{1})\frac{Z_{n}}{2}}\,\rho\,e^{i\theta_{1}\frac{Z_{1}}{2}}e^{i\phi_{2}(\theta_{1})\frac{Z_{2}}{2}}\cdots e^{i\phi_{n}(\theta_{1})\frac{Z_{n}}{2}},$ (SM19)
where $\rho$ is the resource state outputted by the verification procedure. To this aim, we start from the QFI formula in (SM14) and derive a generalized version of (SM16) which is valid when the state is encoded with multiple parameters via unitary encoding:
$\rho_{\theta}=e^{-i\theta A_{1}}e^{-i\phi_{2}(\theta)A_{2}}\cdots e^{-i\phi_{n}(\theta)A_{n}}\,\rho\,e^{i\theta A_{1}}e^{i\phi_{2}(\theta)A_{2}}\cdots e^{i\phi_{n}(\theta)A_{n}},$ (SM20)
with the additional assumption that the generators $A_{j}$ commute pairwise. Then, the encoded resource state in (SM19) is a particular case of (SM20). We start by computing the derivative of the state in (SM20) with respect to $\theta$:
$\partial_{\theta}\rho_{\theta}=U_{\theta}[-iA_{1},\rho]U^{\dagger}_{\theta}+\sum_{k=2}^{n}U_{\theta}[-iA_{k}\phi^{\prime}_{k},\rho]U^{\dagger}_{\theta},$ (SM21)
where we introduced a short-hand notation for the unitary encoding:
$U_{\theta}:=e^{-i\theta A_{1}}e^{-i\phi_{2}(\theta)A_{2}}\cdots e^{-i\phi_{n}(\theta)A_{n}}.$ (SM22)
Then, we observe that the spectral decomposition of $\rho_{\theta}$, in the case of unitary encoding, is obtained from that of $\rho$; namely, the eigenvalues are unchanged ($\lambda_{l}$) and the eigenvectors of $\rho_{\theta}$ are given by: |$\ket{\psi_{l}}=U_{\theta}\ket{\varphi_{l}}$, with |$\ket{\varphi_{l}}$ the eigenvectors of $\rho$. This implies that:
$\braket{\psi_{m}}{\partial_{\theta}\rho_{\theta}}{\psi_{l}}$ $=\bra{\varphi_{m}}\left([-iA_{1},\rho]+\sum_{k=2}^{n}[-iA_{k}\phi^{\prime}_{k},\rho]\right)\ket{\varphi_{l}}$
$=i(\lambda_{m}-\lambda_{l})\left(\braket{\varphi_{m}}{A_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{A_{k}}{\varphi_{l}}\right),$ (SM23)
which substituted in (SM14) yields:
$F_{Q|\theta}(\rho_{\theta})=2\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{A_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{A_{j}}{\varphi_{l}}\right\rvert^{2},$ (SM24)
which is the QFI of the state (SM20), with commuting generators and where $\lambda_{l}$ and |$\ket{\varphi_{l}}$ are obtained from the spectral decomposition of the unencoded state: $\rho=\sum_{l}\lambda_{l}\ket{\varphi_{l}}\bra{\varphi_{l}}$ |. Now, we apply the last expression for the state in (SM19) to solve the optimization problem in (SM18). We obtain:
$\alpha_{1}=\frac{1}{2}\min_{\phi_{k}(\theta),\,k\neq 1}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2},$ (SM25)
where the solution of the optimization, $\alpha_{1}$, can potentially depend on $\theta$ if the derivatives of the functions $\phi_{k}(\theta)$ are not constant. However, we can assume the functions $\phi_{k}(\theta)$ to be of the form: $\phi_{k}(\theta)=c_{k}\theta$, for $c_{k}\in\mathbbm{R}$, without loss of generality. Indeed, suppose that $\tilde{\phi}_{k}(\theta)$ are the optimal functions that solve the optimization problem, and suppose that $\alpha_{1}(\theta)$ depends on $\theta$ through $\tilde{\phi}^{\prime}_{k}(\theta)$. Let $\bar{\theta}$ be the parameter that minimizes $\alpha_{1}$, i.e.: $\bar{\theta}=\arg\min_{\theta}\alpha_{1}(\theta)$. Then, we can choose another set of functions directly proportional to $\theta$, namely $\phi_{k}(\theta)=\theta\tilde{\phi}^{\prime}_{k}(\bar{\theta})$, such that the resulting QFI is minimal and equal to $\alpha_{1}(\bar{\theta})$, with the added benefit of being independent of $\theta$. Therefore, we can restrict the optimization in (SM25) to linear functions of $\theta$. Thus, we get the following parameter for the privacy of $\theta_{1}$ when encoded in the resource state:
$\varepsilon_{1}=\frac{1}{2n^{2}}\min_{c_{k}\in\mathbbm{R}}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{1}}{\varphi_{l}}+\sum_{k=2}^{n}c_{k}\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2}.$ (SM26)
By repeating this procedure for each $j$ and by taking the maximum of the privacy parameters, we obtain the desired $\varepsilon_{p}$ parameter for our resource state.
Hence, the resource state $\rho$, with spectral decomposition $\rho=\sum_{l}\lambda_{l}\ket{\varphi_{l}}\bra{\varphi_{l}}$ |, is $\varepsilon_{p}$-private (according to Definition 1) with respect to the parameters $\theta_{1},\dots,\theta_{n}$ when encoded by $U_{j}=e^{-i\theta_{j}\frac{{Z_{j}}}{2}}$, where:
$\varepsilon_{p}=\frac{1}{2n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{j}}{\varphi_{l}}+\sum_{k\neq j}c_{k|j}\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2}.$ (SM27)
Remark: We verify that the GHZ state encoded by $U_{j}=e^{-i\theta_{j}\frac{{Z_{j}}}{2}}$, for $j=1,\dots,n$, is 00-private. A simple argument follows directly from Definition 1. Indeed, one can choose $\phi_{k}(\theta_{j})=-\theta_{j}/(n-1)$, for $k\neq j$, and observe that the resulting encoded state is independent of $\theta_{j}$:
$\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}}=U_{1}U_{2}\cdots U_{n}\ket{GHZ}\bra{GHZ}U^{\dagger}_{1}U^{\dagger}_{2}\cdots U^{\dagger}_{n}=\ket{GHZ}\bra{GHZ}.$ | . (SM28)
Thus, by definition of the QFI in (SM14), it holds: $F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})=0$ for every $j$ and hence $\varepsilon_{p}=0$ according to Definition 1. Alternatively, we can directly evaluate the privacy parameter from (SM27). For this, we start from the spectral decomposition of the GHZ state (where we identified |$\ket{\varphi_{1}}=\ket{GHZ}$):
[…]
which employed in (SM27) yields:
$\varepsilon_{p}$ $=\frac{1}{n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\sum_{m>1}\left\lvert\braket{\varphi_{m}}{Z_{j}}{GHZ}+\sum_{k\neq j}c_{k|j}\braket{\varphi_{m}}{Z_{k}}{GHZ}\right\rvert^{2}$
$=\frac{1}{n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\sum_{m>1}\left\lvert\bra{\varphi_{m}}\left(Z_{j}+\sum_{k\neq j}c_{k|j}Z_{k}\right)\ket{GHZ}\right\rvert^{2}$
$=\frac{1}{n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\sum_{m>1}\bra{GHZ}\left(Z_{j}+\sum_{k\neq j}c_{k|j}Z_{k}\right)^{\dagger}\ket{\varphi_{m}}\bra{\varphi_{m}}\left(Z_{j}+\sum_{k\neq j}c_{k|j}Z_{k}\right)\ket{GHZ}$
$=\frac{1}{n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\bra{GHZ}\left(Z_{j}+\sum_{k\neq j}c_{k|j}Z_{k}\right)^{\dagger}\left(\mathbbm{1}-\ket{GHZ}\bra{GHZ}\right)\left(Z_{j}+\sum_{k\neq j}c_{k|j}Z_{k}\right)\ket{GHZ}$
$=\frac{1}{n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\left[\braket{GHZ}{(Z^{\prime}_{j})^{2}}{GHZ}-\left(\braket{GHZ}{Z^{\prime}_{j}}{GHZ}\right)^{2}\right],$ (SM30)
where we defined:
$Z^{\prime}_{j}=Z_{j}+\sum_{k\neq j}c_{k|j}Z_{k}.$ (SM31)
Now, due to the structure of the GHZ state, we can e.g. choose $c_{k|j}=-1/(n-1)$, such that $Z^{\prime}_{j}\ket{GHZ}=0$, for every $j$. This yields $\varepsilon_{p}=0$, as claimed.